OpenStudy (anonymous):

3/x+4-7=-4/x+4 (x not equal to -4)

OpenStudy (anonymous):

\[\frac{ 3 }{ x+4 }-7=\frac{ -4 }{ x+4 } (x \neq-4)\]

OpenStudy (anonymous):

i came up with x=-3